# Improving generalisation via anchor multivariate analysis

### v2

@@ 3 Robustness through anchor regularisation
Given $X\in\mathcal{X}$ and $Y\in\mathcal{Y}$, distributionally robust estimation can be formulated as finding a minimiser of the supremum of the population loss under a certain class of distributions:
where $\mathbb{Q}$ is a class of distributions and $\mathcal{L}(X,Y;\mathbf{\Theta})$ is a loss function over $X$ and $Y$ with parameters $\mathbf{\Theta}$. In essence, we aim for our model to perform well even in the worst-case scenario when the distribution of $(X,Y)$ belongs to the class of $\mathbb{Q}$. The choice of the class $\mathbb{Q}$ is particularly important as different definitions lead to various types of robustness.
The class $\mathbb{Q}$, often referred to as the uncertainty set, can be constrained using methods such as $f$-divergence (see, e.g., Namkoong and Duchi, 2016) or the Wasserstein distance (see, e.g., Esfahani and Kuhn, 2015). The challenge of OOD generalisation is also closely related to the field of transfer learning (see, e.g., Zhuang et al., 2021). For an in-depth review of OOD generalisation, we recommend consulting (Shen et al., 2021; Zhou et al., 2023).
Of particular interest for this work, causal inference could be understood as a distributionally robust estimation where the class $\mathbb{Q}$ is the class of all distributions arising from arbitrarily large interventions on the variables of an SCM (see Shen et al., 2021, section 4.1.1).
Besides, Rothenhäusler et al. (2018) proposed to bound the intervention strength on the exogenous variable and relax the assumption that it only affects $Y$ through $X$, considering the linear SCM in Fig. 1. By considering the set of distributions $C^{\gamma}$ defined in (12), they demonstrate that their estimator is optimal in the sense of (14) for the $\ell_{2}$ loss.
@@
The distributional robustness properties stemming from intervention on the anchor variables arise from the ability to bound the $\Sigma^{do(A:=\nu)}_{XY}$ using the observational covariance matrix $\Sigma_{XY}$ and $\Sigma_{XY|A}$. Consequently, any loss function expressed as a linear function of $\Sigma_{XY}$ can attain distributional robustness similar to AR using anchor regularisation.
@@
We say that a loss function $\mathcal{L}(X,Y;\mathbf{\Theta})$ is anchor-compatible if it can be written as
$\mathcal{L}(X,Y;\mathbf{\Theta})=f_{\mathbf{\Theta}}(C_{XY}),$ (15)
where $f_{\mathbf{\Theta}}:\mathbb{R}^{d\times p}\to\mathbb{R}$ is a linear form and $\mathbf{\Theta}$ are the parameters to be learned.
This theorem extends Th. 1 in Rothenhäusler et al. (2018) to the previously defined class of loss functions.
@@
$\begin{split}\sup_{\nu\in C^{\gamma}}&\operatorname{\mathbb{E}_{\nu}}[\mathcal{L}(X,Y;\mathbf{\Theta})]=f_{\mathbf{\Theta}}(\Sigma_{XY})\\ +&(\gamma-1)f_{\mathbf{\Theta}}(\Sigma_{XY|A}),\end{split}$ (16)
where $C^{\gamma}=\{\operatorname{\mathbb{P}_{\nu}}:\Sigma_{\nu}\preceq\gamma\Sigma_{A}\}$.
Proof can be found in the appendix (3.2).
Let us note that the term $f_{\mathbf{\Theta}}(\Sigma_{XY|A})$ could be understood as a causal regulariser enforcing invariance of the loss w.r.t the anchor variable.
For $\gamma=1$, anchor regularisation aligns with unregularised estimation. The theorem suggests that in the case of distributional shifts observed during testing due to constrained intervention on $A$, the anchor-regularised estimator
$\mathbf{\Theta}^{\gamma}=\operatorname{\text{argmin}}_{\mathbf{\Theta}}f_{\mathbf{\Theta}}(\Sigma_{XY})+(\gamma-1)f_{\mathbf{\Theta}}(\Sigma_{XY|A})$ (17)
acts as optimal estimators, as per the definition provided in Eq. (14). We will proceed to demonstrate the practical application of this general result to standard Multivariate Analysis (MVA) algorithms.
@@ 3.2 Common Multivariate Analysis Algorithms
Loss $\|Y-\mathbf{W}^{T}X\|_{F}^{2}$ $\|Y-\mathbf{U}\mathbf{V}^{T}X\|_{F}^{2}$ $\|Y-\mathbf{W}X\|_{F}^{2}$ $\text{tr}\left(\mathbf{W}_{x}^{T}X^{T}Y\mathbf{W}_{y}\right)$ $\text{tr}\left(\mathbf{W}_{x}^{T}X^{T}Y\mathbf{W}_{y}\right)$
@@
A common and straightforward approach to extend least squares (LS) regression to multivariate settings is treating responses independently in multioutput regression (Borchani et al., 2015). This approach aims to identify the regression coefficients $\mathbf{W}$ such that $\hat{Y}=\mathbf{W}^{T}X$, e.g. by solving the LS problem $\widehat{\mathbf{W}}=\operatorname{\text{argmin}}_{\mathbf{W}}\|Y-\mathbf{W}^{T}X\|^{2}_{F}$. When the columns of $Y$ exhibit correlation, RRR addresses a similar optimisation problem assuming that the rank of $\mathbf{W}$ is lower than $\min(d,p)$, see Izenman (1975). Alternatively, in OPLS Arenas-García and Gómez-Verdejo (2015), we do not assume that $\mathbf{W}$ is low rank. Instead, $\mathbf{W}$ is determined by imposing the constraints $\mathbf{W}=\mathbf{V}^{T}\mathbf{U}$, where both $\mathbf{U}$ and $\mathbf{V}$ have a rank $\rho\leq\min(d,p)$.
Additionally, OPLS is commonly solved through matrix deflation, eigenvalue, or generalised eigenvalue decomposition Arenas-García and Gómez-Verdejo (2015). An advantageous property of OPLS is that the columns of the solution $\mathbf{V}$ are ordered based on their predictive performance on $Y$. This is valuable when predicting $Y$ is not the sole objective, and there is interest in learning a relevant information-retaining subspace of $X$ (here $\mathbf{V}^{T}X$). These are natural extensions of LS regression and are anchor-compatible. Please refer to Appendix 8 for further details and proofs.
PLS and CCA are typically preferred when prioritizing learning a pertinent subspace of $X$ over predictive performance. These algoritms aim to maximise the similarity between the latent representations of predictor and target variables. Specifically, PLS assumes the following latent representation:
$X$ $=\mathbf{P}^{T}T+N_{x},$
$Y$ $=\mathbf{Q}^{T}U+N_{y},$
and seeks to maximise the covariance between $\mathbf{W}_{x}^{T}X$ and $\mathbf{W}_{y}^{T}Y$, while CCA aims to maximise the correlation between the estimated latent spaces. Correlation as a similarity measure ensures equal importance for each dimension in the learned latent space, independent of data variance.
PLS, being anchor (see Proposition 8.2), benefits from causal regularisation for distributional robustness. In contrast, CCA lacks anchor compatibility due to nonlinearity in its loss function concerning variance-covariance (see §8.3). As illustrated in Fig. 2, CCA shows reduced robustness under interventions on the anchor variable’s variance. Recognizing the established equivalence between CCA and OPLS (Sun et al., 2009), it may appear puzzling that one algorithm is anchor-compatible while the other is not. We delve deeper into this matter in the supplementary (§8).
4 Towards causality, invariance and stability
Given the well-established connection between causality and invariant predictions (see, for instance Bühlmann (2020) Prop. 3.1), exploring the relationship between anchor regularisation, causality, and invariance could offer valuable insights.
For instance, IV regression assumes that the noise term $\varepsilon_{y}$ is independent of the instrument (or anchor) variable $A$ (i.e., $\varepsilon_{y}\perp\!\!\!\perp A)$ thus IV estimations can be performed using the moment restriction Newey and Powell (2003) $\mathbb{E}[Y-b^{T}X|A]=0$, which can be related to the right term of Eq. (16) for anchor-regularised $\ell_{2}$ loss. [^3] The right term is exactly $\mathbb{E}[(\mathbb{E}[Y-b^{T}X|A])^{2}]$ in AR. Thus, as $\gamma\to\infty$, the anchor-regularised loss converges to the IV estimation, i.e.
@@ 4 Towards causality, invariance and stability
$\mathbf{\Theta}^{IV}$ $=\operatorname{\text{argmin}}_{\mathbf{\Theta}}f_{\mathbf{\Theta}}(\Sigma_{XY|A})=\lim_{\gamma\to+\infty}\mathbf{\Theta}^{\gamma}.$ (18)
Increasing $\gamma$ ensures stable predictive performance by promoting robustness to shifts in the distribution of $X$ resulting from interventions on the anchor.
Another notable scenario is the partialling-out (PA) setting, closely related to double machine learning (see e.g. Chernozhukov et al., 2017), where the anchor variables directly influence both $X$ and $Y$, and there is no hidden confounder $H$.
In this setup, the causal effect is estimated by deriving residuals of $X$ and $Y$ predicted by $A$, followed by estimating the regression coefficient of the residuals of $X$ on the residuals of $Y$. This process is equivalent to estimating the causal effect of $X$ on $Y$ within the orthogonal complement space of $A$. Thus, an anchor-regularised algorithm with $\gamma=0$ is akin to PA estimation.
@@
$\mathbf{\Theta}^{PA}$ $=\operatorname{\text{argmin}}_{\mathbf{\Theta}}f_{\mathbf{\Theta}}(\Sigma_{XY|A^{\perp}})=\lim_{\gamma\to 0}\mathbf{\Theta}^{\gamma},$ (19)
where $A^{\perp}=\mathbf{I}-A$ is the orthogonal complement of $A$. Hence, PA solves the minimax problem Eq. (16) for an empty set Eq. (12) and is thus optimal under distribution shift arising from null intervention on $A$. Note that, in some instances, the parameters $\mathbf{\Theta}^{\gamma}$ remain consistent across all $\gamma\in[0,\infty]$, resulting in what we term as anchor stability. Under these circumstances, and subject to relatively mild assumptions outlined in Rothenhäusler et al. (2018) section 3.2, the parameters $\mathbf{\Theta}$ can be causally interpreted, as demonstrated in (Rothenhäusler et al., 2018, Th. 4), indicating the absence of a confounder $H$ between $X$ and $Y$.
@@ 5 Estimators
In the previous sections, we derived properties of various MVA methods in the population case. We consider now the sample setting where we are given $n$ i.i.d observations of $(X,Y,A)$. As customary, observations are collectively organised row-wise, in the following matrices, $\mathbf{X}\in\mathbb{R}^{n\times d}$, $\mathbf{Y}\in\mathbb{R}^{n\times p}$ and $\mathbf{A}\in\mathbb{R}^{n\times q}$. For a given anchor-compatible loss $\mathcal{L}(X,Y;\mathbf{\Theta})=f_{\mathbf{\Theta}}(C_{XY})$, we propose to consider a simple estimator of the population parameter $\mathbf{\Theta}^{\gamma}$, obtained by the plug-in principle:

### v3

@@ 3 ROBUSTNESS THROUGH ANCHOR REGULARISATION
Building on the aforementioned idea, we demonstrate that anchor regularisation can be efficiently applied to a broad class of linear algorithms, including several classical MVA algorithms. This approach ensures distributional robustness for algorithms within a class $\mathbb{Q}$, which encompasses distributions arising from bounded interventions on $A$.
@@ 3.1 Anchor-Compatible Loss
The distributional robustness properties stemming from intervention on the anchor variables arise from the ability to bound $\Sigma^{do(A\sim\nu)}_{XY}$ using the observational covariance matrices $\Sigma_{XY}$ and $\Sigma_{XY|A}$. Consequently, any loss function expressed as a linear form on $\Sigma_{XY}$ can attain distributional robustness similar to AR using anchor regularisation.
@@
We say that a loss function $\mathcal{L}(X,Y;\mathbf{\Theta})$ is anchor-compatible if it can be written as $\mathcal{L}(X,Y;\mathbf{\Theta})=f_{\mathbf{\Theta}}(C_{XY})$, where $f_{\mathbf{\Theta}}:\mathbb{R}^{d\times p}\to\mathbb{R}$ is a linear form [^3] A linear form is a linear map from a vector space to its field of scalars. The trace function is, for example, a linear form. and $\mathbf{\Theta}$ are the parameters to be learned.
The following result extends Th. 1 in Rothenhäusler et al. (2018) to the previously defined class of loss functions.
@@
$\begin{split}\sup_{\nu\in C^{\gamma}}\operatorname{\mathbb{E}_{\nu}}[\mathcal{L}(X,Y;\mathbf{\Theta})]&=f_{\mathbf{\Theta}}(\Sigma_{XY})\\ &+(\gamma-1)f_{\mathbf{\Theta}}(\Sigma_{XY|A}),\end{split}$ (7)
where $C^{\gamma}=\{\operatorname{\mathbb{P}_{\nu}}:\Sigma_{\nu}\preceq\gamma\Sigma_{A}\}$. Proof can be found in Appendix (3.2).
The term $f_{\mathbf{\Theta}}(\Sigma_{XY|A})$ could be understood as a causal regulariser enforcing invariance of the loss w.r.t the anchor variable. More intuitively, the anchor regularisation can be seen as adding a penalty term to how much $X$ and $Y$ covary when projected in the span of the anchor variable. The theorem suggests that in the case of distributional shifts observed during testing due to constrained intervention on $A$, the anchor-regularised estimator $\mathbf{\Theta}^{\gamma}=\operatorname{\text{argmin}}_{\mathbf{\Theta}}f_{\mathbf{\Theta}}(\Sigma_{XY})+(\gamma-1)f_{\mathbf{\Theta}}(\Sigma_{XY|A})$ acts as an optimal estimator, as per the definition provided in Eq. (1). We will proceed to demonstrate the practical application of this general result to standard MVA algorithms.
@@ 3.2 Common Multivariate Analysis Algorithms
Loss $\|Y-\mathbf{W}^{T}X\|_{F}^{2}$ $\|Y-\mathbf{U}\mathbf{V}^{T}X\|_{F}^{2}$ $\|Y-\mathbf{W}X\|_{F}^{2}$ $-\text{tr}\left(\mathbf{W_{x}}^{T}X^{T}Y\mathbf{W_{y}}\right)$ $-\text{tr}\left(\mathbf{W_{x}}^{T}X^{T}Y\mathbf{W_{y}}\right)$
@@
$\mathbf{W_{y}}^{T}\mathbf{W_{y}}=\mathbf{I}$ $\mathbf{W_{y}}^{T}C_{Y}\mathbf{W_{y}}=\mathbf{I}$
@@
By virtue of Th. 3.2, any anchor-compatible loss function has theoretically grounded robustness properties to distribution shifts. This is practically relevant, as users can easily verify if a loss function is compatible using Def. 3.1. In that section, we show how this can be applied to a set of commonly used multivariate algorithms in both compatible and incompatible cases. As demonstrated in the experimental section, while non-compatible loss functions might also benefit from anchor regularisation, they exhibit suboptimal robustness properties, and this regularisation should thus be used with care.
A common and straightforward approach to extend Least Squares (LS) regression to multivariate settings is treating responses independently in multioutput regression (Borchani et al., 2015). This approach aims to identify the regression coefficients $\mathbf{W}$ such that $\hat{Y}=\mathbf{W}^{T}X$, e.g. by solving the LS problem $\widehat{\mathbf{W}}=\operatorname{\text{argmin}}_{\mathbf{W}}\|Y-\mathbf{W}^{T}X\|^{2}_{F}$. When the columns of $Y$ exhibit correlation, RRR addresses a similar optimisation problem assuming that the rank of $\mathbf{W}$ is lower than $\min(d,p)$, see Izenman (1975). Alternatively, in OPLS (Arenas-García and Gómez-Verdejo, 2015), we do not assume that $\mathbf{W}$ is low rank. Instead, $\mathbf{W}$ is determined by imposing the constraints $\mathbf{W}=\mathbf{V}^{T}\mathbf{U}$, where both $\mathbf{U}$ and $\mathbf{V}$ have a rank $\rho\leq\min(d,p)$. Additionally, OPLS is commonly solved through matrix deflation, eigenvalue, or generalised eigenvalue decomposition (Arenas-García and Gómez-Verdejo, 2015). An advantageous property of OPLS is that the columns of the solution $\mathbf{V}$ are ordered based on their predictive performance on $Y$. This is valuable when predicting $Y$ is not the sole objective, and there is interest in learning a relevant information-retaining subspace of $X$ (here $\mathbf{V}^{T}X$). These are natural extensions of LS regression and are anchor-compatible. See Appendix A for further details and proofs. PLS and CCA are typically preferred when prioritizing learning a pertinent subspace of $X$ over predictive performance. These algoritms aim to maximise the similarity between the latent representations of predictor and target variables. Specifically, PLS assumes the following latent representation $X=\mathbf{P}^{T}T+N_{x}$ and $Y=\mathbf{Q}^{T}U+N_{y}$ and seeks to maximise the covariance between $\mathbf{W}_{x}^{T}X$ and $\mathbf{W}_{y}^{T}Y$, while CCA aims to maximise the correlation between the estimated latent spaces. Correlation as a similarity measure ensures equal importance for each dimension in the learned latent space, independent of data variance. PLS, being anchor-compatible (see Prop. A.2), benefits from causal regularisation for distributional robustness. In contrast, CCA lacks anchor compatibility due to nonlinearity in its loss function concerning variance-covariance (see §A.3). As illustrated in Fig. 2, CCA shows reduced robustness under interventions on the anchor variable’s variance. Recognizing the established equivalence between CCA and OPLS (Sun et al., 2009), it may appear puzzling that one algorithm is anchor-compatible while the other is not. We delve deeper into this matter in the Appendix (see §A).
4 TOWARDS INVARIANCE AND CAUSALITY
A direct implication of Th. 3.2 is the connection between anchor regularisation and two well-studied estimators: Instrumental Variable (IV) and Partialling-Out (PA), which we define as follow
@@ 4 TOWARDS INVARIANCE AND CAUSALITY
$\mathbf{\Theta}^{IV}$ $=\operatorname{\text{argmin}}_{\mathbf{\Theta}}f_{\mathbf{\Theta}}(\Sigma_{XY|A})=\lim_{\gamma\to+\infty}\mathbf{\Theta}^{\gamma}$ (8)
$\mathbf{\Theta}^{PA}$ $=\operatorname{\text{argmin}}_{\mathbf{\Theta}}f_{\mathbf{\Theta}}(\Sigma_{XY|A}^{\perp})=\mathbf{\Theta}^{0}.$ (9)
Here $\Sigma_{XY|A}^{\perp}=\Sigma_{XY}-\Sigma_{XY|A}$. Note that for the MLR loss function $\mathbf{\Theta}^{IV}$ is equivalent to the two-stage least square estimation. A direct implication of Th. 3.2 is that the minimisers of the anchor-regularised loss function have the following properties: unregularised parameters $\mathbf{\Theta}^{1}$ are optimal for all $\nu\in C^{1}$, PA parameters $\mathbf{\Theta}^{PA}$ are optimal for all $\nu\in C^{0}$ and IV parameters $\mathbf{\Theta}^{IV}$ are optimal for all $\nu\in C^{\infty}$. However, although PA and IV estimations are generally used for causal inference, this shows that they provide robustness properties when considering specific interventions on the anchor variables, even if they do not necessarily retrieve the causal parameters. While it is known that for specific sets of interventions, causal parameters have robustness properties (Haavelmo, 1943), it has recently been recognized (Christiansen et al., 2022) that this is not always the case. The following proposition shows the connection between the optimal structural parameters in terms of error and the causal parameters.
Proposition 4.1.
Let us denote $R(\tilde{\mathbf{B}})=[X,Y,H]^{T}-\tilde{\mathbf{B}}[X,Y,H]^{T}$ as the error term when reconstructing $(X,Y,H)$ with $\tilde{\mathbf{B}}$. Given any matrix $\operatorname{\mathbf{B}^{\diamond}}$ such that
$\begin{split}\sup_{\nu\in C^{\gamma}}\operatorname{\mathbb{E}_{\nu}}[f_{\mathbf{\Theta}}(R(\operatorname{\mathbf{B}^{\diamond}})&\otimes R(\operatorname{\mathbf{B}^{\diamond}}))]\\ &\leq\sup_{\nu\in C^{\gamma}}\operatorname{\mathbb{E}_{\nu}}[f_{\mathbf{\Theta}}(R(\mathbf{B})\otimes R(\mathbf{B}))],\end{split}$
it holds that if the loss $f_{\mathbf{\Theta}}$ is anchor-compatible, then
@@
$\begin{split}\sup_{\nu\in C^{\gamma}}f_{\mathbf{\Theta}}(&(\mathbf{B}-\operatorname{\mathbf{B}^{\diamond}})\Sigma_{XYH}^{do(A\sim\nu)}(\mathbf{B}-\operatorname{\mathbf{B}^{\diamond}})^{T})\\ &\leq 4f_{\mathbf{\Theta}}(\Sigma_{\epsilon})+4\gamma f_{\mathbf{\Theta}}(\mathbf{M}\Sigma_{A}\mathbf{M}^{T}).\end{split}$ (10)
The worst case risk between the causal parameter $\mathbf{B}$ and the optimal one $\operatorname{\mathbf{B}^{\diamond}}$ is bounded but increase linearly with $\gamma$. Thus the potential improvement enabled by using anchor-regularisation instead of seeking to retrieve causal parameters becomes clearer as intervention strength on the anchor increases.
@@ 5 ESTIMATORS
In the previous sections, we derived properties of a regularised version of various MVA methods in the population case. We consider now the sample setting where we are given $n$ i.i.d observations of $(X,Y,A)$. Observations are collectively organised row-wise, in the following matrices, $\mathbf{X}\in\mathbb{R}^{n\times d}$, $\mathbf{Y}\in\mathbb{R}^{n\times p}$ and $\mathbf{A}\in\mathbb{R}^{n\times q}$. For a given anchor-compatible loss $\mathcal{L}(X,Y;\mathbf{\Theta})=f_{\mathbf{\Theta}}(C_{XY})$, we propose a simple estimator of the population parameters: